# Multiple Proposer Transaction Fee Mechanism Design: Robust Incentives Against Censorship and Bribery

### v1

@@ 3.2 Phases and Allocation Rules
When the type of the includer and the block producer is implied by the context we write $F^{Init}_{j}$ and $F^{Init}_{BP}$
Phase $2$
@@ 3.3 Payments and Costs
{mdframed} Payments when the block is approved (not ignored) by the attesters: Assume that the types of the parties are implied by the context.
@@
{mdframed} Costs when the block is approved by the attesters: 1. The sender of the transaction pays an amount that is burnt by the system (for instance, in Ethereum, this is the base fee, a protocol-computed reserve price determined by the burning rule of EIP-1559 [11]).
@@
{mdframed} Costs regardless whether the block is approved by the attesters: Assume that the types of the parties are implied by the context. 1. The block proposer pays for every fake transaction $t$ they have included in the mempool during Phase $1$ ($t\in F^{Init}_{BP}$) $\gamma$ fraction of the payment awarded to the block producer and the committee for this transaction ($p^{BP}_{t}(H,B_{k},\alpha)$ and $\sum_{j=1}^{m}p^{CM}_{t}(H,B_{k},\alpha,j)$ respectively, multiplied by $s_{t}$). This captures the scenario where the block producer is unsure whether they will be the creator of $B_{k}$ so that they can invalidate their initial fake transactions. As a result, they risk forfeiting the associated fees, which would otherwise go to the creator of $B_{k}$ and the includers. If the block producer is certain they will be the creator of $B_{k}$, then $\gamma=0$. 2. If the includers and the block producer include a transaction initiated by the users in their inclusion list or block respectively, they forfeit the reward they would have received from the briber if they had excluded it.
@@ 3.4 Utilities
{mdframed} Utility of the user when the types of the parties are fixed and known: At a high level, when the transaction of the user is included in the current block and the block is approved by the attesters, they gain the value of the transaction per unit of size. Moreover, they pay the priority fee to the includers and the block producers (if any, depending on the payment mechanism), and the burning fee, per unit of size.
@@
{mdframed} Utility of includer $j$ when the types of the parties are fixed and known: At a high level, if $B_{k}$ is approved by the attesters, for every transaction in $B_{k}$ not originated from the includer (this means that it does not belong to $F_{j}\cup F^{Init}_{j}$), the includer receives the corresponding transaction fee. In addition, for every transaction from $M_{0}$ in their inclusion list, they lose the bribe. Furthermore, for every transaction in their inclusion list, not created by them, they lose the per unit of size cost $\mu_{CM}^{Cost}$ Finally, if $B_{k}$ is approved by the attesters, they pay the burning fee and the payment for the block producer and the other includers, per unit of size for every fake transaction in $B_{k}$ they have submitted to the mempool or they have included in their inclusion list.
@@
Utility of the includer $j$. If $B_{k}$ meets protocol’s predetermined criteria for inclusion lists: $u_{CM}(b_{t},H,B_{k},\alpha,j)=$ $\sum_{t\in B_{k}\cap\neg(F_{j}\cup F^{Init}_{j})}[p^{CM}_{t}(H,B_{k},\alpha,j)\cdot s_{t}]$ $-\sum_{t\in\mathsf{IL_{j}}\cap M_{0}}[\mu_{CM_{j}}^{Bribe}(s_{t},b_{t},p_{t},\vec{\alpha}_{t},H,M_{0})]-\sum_{t\in\mathsf{IL_{j}}\cap\neg(F_{j}\cup F^{Init}_{j})}[\mu_{CM}^{Cost}\cdot s_{t}]$ $-\sum_{t\in B_{k}\cap(F_{j}\cup F^{Init}_{j})}[q_{t}(H,B_{k})+p^{BP}_{t}(H,B_{k},\alpha)+\sum_{i\in\{1,\dots,m\}\setminus j}p^{CM}_{t}(H,B_{k},\alpha,i)]\cdot s_{t}$ Otherwise: The same as above with the difference that the first and the third term are equal to 00.
{mdframed} Utility of the producer when the types of the parties are fixed and known: At a high level, if $B_{k}$ is approved by the attesters, for every transaction in $B_{k}$, the block producer receives the corresponding fee and incurs the cost $\mu_{BP}^{Cost}$, per unit of size (except for the transactions that originated from the block producer). Moreover, regardless whether $B_{k}$ is approved by the attesters, for every transaction in $M_{0}$, the block producer loses the bribe. Additionaly, if $B_{k}$ is approved by the attesters, for every fake transaction in $B_{k}$ created by them, they pay the burning fee and the fee for the includers per unit of size. Finally, for every transaction they submitted to the mempool in Phase $1$, they pay $\gamma$ fraction of the fee that corresponds to the block producer and the committee.
@@
Utility of the block producer. Let $Fee_{t}$ be the total fee per unit of size that a transaction $t$ gives to the committee and the block producer when it is included in both an inclusion list and a block. The structure of this fee depends on the specifics of the bid $b_{t}$ and the fee mechanism under examination. If $B_{k}$ meets protocol’s predetermined criteria for inclusion lists: $u_{BP}(b_{t},H,B_{k},\alpha)$ $=\sum_{t\in B_{k}\cap\neg(F_{BP}\cup F^{Init}_{BP})}[(p^{BP}_{t}(H,B_{k},\alpha)-\mu_{BP}^{Cost})\cdot s_{t}]$ $-\sum_{t\in B_{k}\cap M_{0}}[\mu_{BP}^{Bribe}(s_{t},b_{t},p_{t},\vec{\alpha}_{t},H,M_{0})]$ $-\sum_{t\in B_{k}\cap(F^{Init}_{BP}\cup F_{BP})}[(q_{t}(H,B_{k})+\sum_{i\in\{1,\dots,m\}}p^{CM}_{t}(H,B_{k},\alpha,i))\cdot s_{t}]$ $-\sum_{t\in F^{Init}_{BP}}[\gamma\cdot Fee_{t}\cdot s_{t}]$ Otherwise: The same with the difference that the first and the third term are equal to 00.
@@ 3.5 Formal Definition of the Multiple Proposer Transaction Fee Mechanism and its Properties
($(\{x^{BP,\mu_{BP}^{Bribe}}\}_{\mu_{BP}^{Bribe}\in\text{Bribe}^{BP}},\{x^{j,\mu_{CM_{j}}^{Bribe}}\}_{\forall\mu_{CM_{j}}^{Bribe}\in\text{Bribe}^{CM}},p^{CM},p^{BP},q)$
@@ 4 Double TFM
{mdframed} Notation Let $t_{0}$ be the target transaction that the external briber seeks to censor by offering bribes to the includers and/or block producer. Let $c_{Incl}:=\lfloor C_{Incl}/s\rfloor$ be the maximum number of transactions an inclusion list can store. Let $c_{block}:=\lfloor C_{block}/s\rfloor$ be the maximum number of transactions a block can store. Let $r:=q_{t}(H,B_{k})=q_{t}(H)$ be the burning fee per unit of size for a transaction $t$ included in $B_{k}$ according to EIP-1559. Recall that $H$ is the history of blocks. Let $w$ be the number of transactions in $M_{0}$. Given $M_{0}$, we define the following two ordered lists of transactions: $L_{BP}$: This list consists of the transactions in $M_{0}$ that have block producer fee no smaller than $\mu^{Cost}_{BP}\cdot s$. This list is ordered by the block producer fee. The ordering is decreasing. The ties break according to a deterministic rule. The block producer fee corresponding to position $j$ is denoted by $f_{j,BP}$. If there is no position $j$ in the list, we consider that $f_{j,BP}=0$. $L_{CM,c_{block}}$: This list consists of the transactions in $M_{0}$ that (i) belong to the first $c_{block}$ positions in $L_{BP}$ (or to $L_{BP}$ if $L_{BP}$ has fewer than $c_{block}$ positions) (ii) have committee fee no smaller than $\mu^{Cost}_{CM}\cdot s$. This list is ordered by the committee fee. The ordering is decreasing. The ties break according to a deterministic rule. The committee fee corresponding to position $j$ is denoted by $f_{j,CM}$. We assume that $t_{0}$ belongs to the first $\min\{c_{Incl}\cdot m,size_{L_{CM,c_{block}}}\}$ positions in $L_{CM,c_{block}}$, where $size_{L_{CM,c_{block}}}$ is the size of $L_{CM,c_{block}}$. We adopt this assumption because, under the allocation rule defined below, any transaction that fails to meet these requirements would not be included in either the inclusion list or the block, even in the absence of a bribe. Let $size_{L_{BP}}$ be the size of $L_{BP}$. Let $sum_{max,c_{block}}$be the sum of the rewards (block producer fee minus $\mu^{Cost}_{BP}\cdot s$ per transaction) the block producer will receive if they include the first $\min\{c_{block},size_{L_{BP}}\}$ transactions of $L_{BP}$ in their block. Let $o$ be the order of $t_{0}$ in $L_{CM,c_{block}}$ and $o_{BP}$ be the order of $t_{0}$ in $L_{BP}$. Let $f_{CM},f_{BP}$ be the committee and the block producer fee of $t_{0}$ respectively.
@@ 4.1 Formal Definition of Double TFM
{mdframed} Set $\text{Bribe}^{CM}$ The $\text{Bribe}^{CM}$consists of the following three functions. $\mu^{Bribe}_{1,CM}$: This function gives to the includer $f_{CM}-\max\{f_{g,CM},\mu^{Cost}_{CM}\cdot s\}$, where $g=c_{Incl}\cdot\lceil\frac{o}{c_{Incl}}\rceil+1$ if they omit this transaction (regardless of the strategy of the other includers). $\mu^{Bribe}_{2,CM}$: This function gives to the includer $f_{CM}-\max\{f_{g,CM},\mu^{Cost}_{CM}\cdot s\}$, where $g=c_{Incl}\cdot\lceil\frac{o}{c_{Incl}}\rceil+1$ if they omit this transaction and all the other includers and the block producer do the same. $\mu^{Bribe}_{3,CM}$: This function gives an $X$ to the includer if they omit this transaction (regardless of the strategy of the other includers). $X$ can be any non negative real number.
{mdframed} Set $\text{Bribe}^{BP}$ $\text{Bribe}^{BP}$consists of the following bribe function $\mu^{Bribe}_{1,BP}$. When $w>c_{block}$: This function gives to the block producer $f_{BP}-\max\{f_{c_{block}+1,BP},\mu^{Cost}_{BP}\cdot s\}$ if they omit the transaction regardless of the strategy of the includers. When $w\leq c_{block}$. If at most one transaction per sender is allowed to be added to an inclusion list, the function $\mu^{Bribe}_{1,BP}$ gives to the block producer: $\min\{f_{BP}-\mu^{Cost}_{BP}\cdot s+r\cdot s\cdot(c_{block}-w+1),$ $sum_{max,c_{block}},$ $f_{BP}-\mu^{Cost}_{BP}\cdot s$ $+(c_{block}-o_{BP}+1)\cdot\gamma\cdot f_{BP}+\lceil\frac{(c_{block}-o_{BP}+1)}{m}\rceil\cdot r\cdot s,$ $f_{BP}-\mu^{Cost}_{BP}\cdot s+$ $\gamma\cdot(\min\{m\cdot c_{Incl},size_{L_{CM,c_{block}}}\}-o+1)\cdot$ $(f_{CM}+\max\{f_{c_{block}-(\min\{m\cdot c_{Incl},size_{L_{CM,c_{block}}}\}-o+1)+1,BP},\mu^{Cost}_{BP}\cdot s\})$ $+\lceil\frac{(\min\{m\cdot c_{Incl},size_{L_{CM,c_{block}}}\}-o+1)}{m}\rceil\cdot r\cdot s\}$ If multiple transactions per sender are allowed to be added to an inclusion list, the function $\mu^{Bribe}_{1,BP}$ gives to the block producer: $\min\{f_{BP}-\mu^{Cost}_{BP}\cdot s+r\cdot s\cdot(c_{block}-w+1),$ $sum_{max,c_{block}},$ $f_{BP}-\mu^{Cost}_{BP}\cdot s+(c_{block}-o_{BP}+1)\cdot\gamma\cdot f_{BP}+r\cdot s,$ $f_{BP}-\mu^{Cost}_{BP}\cdot s+\gamma\cdot(\min\{m\cdot c_{Incl},size_{L_{CM,c_{block}}}\}-o+1)\cdot$ $(f_{CM}+\max\{f_{c_{block}-(\min\{m\cdot c_{Incl},size_{L_{CM,c_{block}}}\}-o+1)+1,BP},\mu^{Cost}_{BP}\cdot s\})+r\cdot s\}$
@@
($(\{x^{BP,\mu^{Bribe}_{BP}}\}_{\mu^{Bribe}_{BP}\in\text{Bribe}^{BP}},\{x^{j,\mu^{Bribe}_{CM_{j}}}\}_{\mu^{Bribe}_{CM_{j}}\in\text{Bribe}^{CM}},p^{CM}p^{BP},q)$
@@
    • Repeats the same procedure for the includers with order $12,\dots,j-$ 1.

### v2

# Aikaterini-Panagiota Stouka Nethermind Research

@@ 3.2 Phases and Allocation Rules
When the type of the includer and the block producer is implied by the context we write $F^{Init}_{j}$ and $F^{Init}_{BP}$ .
Phase 22
@@ 3.3 Payments and Costs
Payments when the block is approved (not ignored) by the attesters:
Assume that the types of the parties are implied by the context.
@@
Costs when the block is approved by the attesters:
1. The sender of the transaction pays an amount that is burnt by the system (for instance, in Ethereum, this is the base fee, a protocol-computed reserve price determined by the burning rule of EIP-1559 [11]).
@@
Costs regardless whether the block is approved by the attesters:
Assume that the types of the parties are implied by the context.
1. The block proposer pays for every fake transaction $t$ they have included in the mempool during Phase 11 ($t\in F^{Init}_{BP}$) $\gamma$ fraction of the payment awarded to the block producer and the committee for this transaction ($p^{BP}_{t}(H,B_{k},\alpha)$ and $\sum_{j=1}^{m}p^{CM}_{t}(H,B_{k},\alpha,j)$ respectively, multiplied by $s_{t}$). This captures the scenario where the block producer is unsure whether they will be the creator of $B_{k}$ so that they can invalidate their initial fake transactions. As a result, they risk forfeiting the associated fees, which would otherwise go to the creator of $B_{k}$ and the includers. If the block producer is certain they will be the creator of $B_{k}$, then $\gamma=0$. 2. If the includers and the block producer include a transaction initiated by the users in their inclusion list or block respectively, they forfeit the reward they would have received from the briber if they had excluded it.
@@ 3.4 Utilities
Utility of the user when the types of the parties are fixed and known:
At a high level, when the transaction of the user is included in the current block and the block is approved by the attesters, they gain the value of the transaction per unit of size. Moreover, they pay the priority fee to the includers and the block producers (if any, depending on the payment mechanism), and the burning fee, per unit of size.
@@
Utility of includer $j$ when the types of the parties are fixed and known:
At a high level, if $B_{k}$ is approved by the attesters, for every transaction in $B_{k}$ not originated from the includer (this means that it does not belong to $F_{j}\cup F^{Init}_{j}$), the includer receives the corresponding transaction fee. In addition, for every transaction from $M_{0}$ in their inclusion list, they lose the bribe. Furthermore, for every transaction in their inclusion list, not created by them, they lose the per unit of size cost $\mu_{CM}^{Cost}$ . Finally, if $B_{k}$ is approved by the attesters, they pay the burning fee and the payment for the block producer and the other includers, per unit of size for every fake transaction in $B_{k}$ they have submitted to the mempool or they have included in their inclusion list.
@@
Utility of the includer $j$. If $B_{k}$ meets protocol’s predetermined criteria for inclusion lists: $u_{CM}(b_{t},H,B_{k},\alpha,j)=$ $\sum_{t\in B_{k}\cap\neg(F_{j}\cup F^{Init}_{j})}[p^{CM}_{t}(H,B_{k},\alpha,j)\cdot s_{t}]$ $-\sum_{t\in\mathsf{IL_{j}}\cap M_{0}}[\mu_{CM_{j}}^{Bribe}(s_{t},b_{t},p_{t},\vec{\alpha}_{t},H,M_{0})]-\sum_{t\in\mathsf{IL_{j}}\cap\neg(F_{j}\cup F^{Init}_{j})}[\mu_{CM}^{Cost}\cdot s_{t}]$ $-\sum_{t\in B_{k}\cap(F_{j}\cup F^{Init}_{j})}[q_{t}(H,B_{k})+p^{BP}_{t}(H,B_{k},\alpha)+\sum_{i\in\{1,\dots,m\}\setminus j}p^{CM}_{t}(H,B_{k},\alpha,i)]\cdot s_{t}$ Otherwise: The same as above with the difference that the first and the third term are equal to 0.
Utility of the producer when the types of the parties are fixed and known:
At a high level, if $B_{k}$ is approved by the attesters, for every transaction in $B_{k}$, the block producer receives the corresponding fee and incurs the cost $\mu_{BP}^{Cost}$, per unit of size (except for the transactions that originated from the block producer). Moreover, regardless whether $B_{k}$ is approved by the attesters, for every transaction in $M_{0}$, the block producer loses the bribe. Additionaly, if $B_{k}$ is approved by the attesters, for every fake transaction in $B_{k}$ created by them, they pay the burning fee and the fee for the includers per unit of size. Finally, for every transaction they submitted to the mempool in Phase $1$, they pay $\gamma$ fraction of the fee that corresponds to the block producer and the committee.
@@
Utility of the block producer. Let $Fee_{t}$ be the total fee per unit of size that a transaction $t$ gives to the committee and the block producer when it is included in both an inclusion list and a block. The structure of this fee depends on the specifics of the bid $b_{t}$ and the fee mechanism under examination. If $B_{k}$ meets protocol’s predetermined criteria for inclusion lists: $u_{BP}(b_{t},H,B_{k},\alpha)$ $=\sum_{t\in B_{k}\cap\neg(F_{BP}\cup F^{Init}_{BP})}[(p^{BP}_{t}(H,B_{k},\alpha)-\mu_{BP}^{Cost})\cdot s_{t}]$ $-\sum_{t\in B_{k}\cap M_{0}}[\mu_{BP}^{Bribe}(s_{t},b_{t},p_{t},\vec{\alpha}_{t},H,M_{0})]$ $-\sum_{t\in B_{k}\cap(F^{Init}_{BP}\cup F_{BP})}[(q_{t}(H,B_{k})+\sum_{i\in\{1,\dots,m\}}p^{CM}_{t}(H,B_{k},\alpha,i))\cdot s_{t}]$ $-\sum_{t\in F^{Init}_{BP}}[\gamma\cdot Fee_{t}\cdot s_{t}]$ Otherwise: The same with the difference that the first and the third term are equal to 0.
@@ 3.5 Formal Definition of the Multiple Proposer Transaction Fee Mechanism and its Properties
$(\{x^{BP,\mu_{BP}^{Bribe}}\}_{\mu_{BP}^{Bribe}\in\text{Bribe}^{BP}},\{x^{j,\mu_{CM_{j}}^{Bribe}}\}_{\forall\mu_{CM_{j}}^{Bribe}\in\text{Bribe}^{CM}},p^{CM},p^{BP},q)$
@@ 4 Double TFM
Notation Let $t_{0}$ be the target transaction that the external briber seeks to censor by offering bribes to the includers and/or block producer. Let $c_{Incl}:=\lfloor C_{Incl}/s\rfloor$ be the maximum number of transactions an inclusion list can store. Let $c_{block}:=\lfloor C_{block}/s\rfloor$ be the maximum number of transactions a block can store. Let $r:=q_{t}(H,B_{k})=q_{t}(H)$ be the burning fee per unit of size for a transaction $t$ included in $B_{k}$ according to EIP-1559. Recall that $H$ is the history of blocks. Let $w$ be the number of transactions in $M_{0}$. Given $M_{0}$, we define the following two ordered lists of transactions: $L_{BP}$: This list consists of the transactions in $M_{0}$ that have block producer fee no smaller than $\mu^{Cost}_{BP}\cdot s$. This list is ordered by the block producer fee. The ordering is decreasing. The ties break according to a deterministic rule. The block producer fee corresponding to position $j$ is denoted by $f_{j,BP}$. If there is no position $j$ in the list, we consider that $f_{j,BP}=0$. $L_{CM,c_{block}}$: This list consists of the transactions in $M_{0}$ that (i) belong to the first $c_{block}$ positions in $L_{BP}$ (or to $L_{BP}$ if $L_{BP}$ has fewer than $c_{block}$ positions) (ii) have committee fee no smaller than $\mu^{Cost}_{CM}\cdot s$. This list is ordered by the committee fee. The ordering is decreasing. The ties break according to a deterministic rule. The committee fee corresponding to position $j$ is denoted by $f_{j,CM}$. We assume that $t_{0}$ belongs to the first $\min\{c_{Incl}\cdot m,size_{L_{CM,c_{block}}}\}$ positions in $L_{CM,c_{block}}$, where $size_{L_{CM,c_{block}}}$ is the size of $L_{CM,c_{block}}$. We adopt this assumption because, under the allocation rule defined below, any transaction that fails to meet these requirements would not be included in either the inclusion list or the block, even in the absence of a bribe. Let $size_{L_{BP}}$ be the size of $L_{BP}$. Let $sum_{max,c_{block}}$be the sum of the rewards (block producer fee minus $\mu^{Cost}_{BP}\cdot s$ per transaction) the block producer will receive if they include the first $\min\{c_{block},size_{L_{BP}}\}$ transactions of $L_{BP}$ in their block. Let $o$ be the order of $t_{0}$ in $L_{CM,c_{block}}$ and $o_{BP}$ be the order of $t_{0}$ in $L_{BP}$. Let $f_{CM},f_{BP}$ be the committee and the block producer fee of $t_{0}$ respectively.
@@ 4.1 Formal Definition of Double TFM
Set $\text{Bribe}^{CM}$
The $\text{Bribe}^{CM}$consists of the following three functions.
$\mu^{Bribe}_{1,CM}$: This function gives to the includer $f_{CM}-\max\{f_{g,CM},\mu^{Cost}_{CM}\cdot s\}$, where $g=c_{Incl}\cdot\lceil\frac{o}{c_{Incl}}\rceil+1$ if they omit this transaction (regardless of the strategy of the other includers). $\mu^{Bribe}_{2,CM}$: This function gives to the includer $f_{CM}-\max\{f_{g,CM},\mu^{Cost}_{CM}\cdot s\}$, where $g=c_{Incl}\cdot\lceil\frac{o}{c_{Incl}}\rceil+1$ if they omit this transaction and all the other includers and the block producer do the same. $\mu^{Bribe}_{3,CM}$: This function gives an $X$ to the includer if they omit this transaction (regardless of the strategy of the other includers). $X$ can be any non negative real number.
Set $\text{Bribe}^{BP}$
$\text{Bribe}^{BP}$consists of the following bribe function $\mu^{Bribe}_{1,BP}$.
When $w>c_{block}$: This function gives to the block producer $f_{BP}-\max\{f_{c_{block}+1,BP},\mu^{Cost}_{BP}\cdot s\}$ if they omit the transaction regardless of the strategy of the includers. When $w\leq c_{block}$. If at most one transaction per sender is allowed to be added to an inclusion list, the function $\mu^{Bribe}_{1,BP}$ gives to the block producer: $\min\{f_{BP}-\mu^{Cost}_{BP}\cdot s+r\cdot s\cdot(c_{block}-w+1),$ $sum_{max,c_{block}},$ $f_{BP}-\mu^{Cost}_{BP}\cdot s$ $+(c_{block}-o_{BP}+1)\cdot\gamma\cdot f_{BP}+\lceil\frac{(c_{block}-o_{BP}+1)}{m}\rceil\cdot r\cdot s,$ $f_{BP}-\mu^{Cost}_{BP}\cdot s+$ $\gamma\cdot(\min\{m\cdot c_{Incl},size_{L_{CM,c_{block}}}\}-o+1)\cdot$ $(f_{CM}+\max\{f_{c_{block}-(\min\{m\cdot c_{Incl},size_{L_{CM,c_{block}}}\}-o+1)+1,BP},\mu^{Cost}_{BP}\cdot s\})$ $+\lceil\frac{(\min\{m\cdot c_{Incl},size_{L_{CM,c_{block}}}\}-o+1)}{m}\rceil\cdot r\cdot s\}$ If multiple transactions per sender are allowed to be added to an inclusion list, the function $\mu^{Bribe}_{1,BP}$ gives to the block producer: $\min\{f_{BP}-\mu^{Cost}_{BP}\cdot s+r\cdot s\cdot(c_{block}-w+1),$ $sum_{max,c_{block}},$ $f_{BP}-\mu^{Cost}_{BP}\cdot s+(c_{block}-o_{BP}+1)\cdot\gamma\cdot f_{BP}+r\cdot s,$ $f_{BP}-\mu^{Cost}_{BP}\cdot s+\gamma\cdot(\min\{m\cdot c_{Incl},size_{L_{CM,c_{block}}}\}-o+1)\cdot$ $(f_{CM}+\max\{f_{c_{block}-(\min\{m\cdot c_{Incl},size_{L_{CM,c_{block}}}\}-o+1)+1,BP},\mu^{Cost}_{BP}\cdot s\})+r\cdot s\}$
@@
$(\{x^{BP,\mu^{Bribe}_{BP}}\}_{\mu^{Bribe}_{BP}\in\text{Bribe}^{BP}},\{x^{j,\mu^{Bribe}_{CM_{j}}}\}_{\mu^{Bribe}_{CM_{j}}\in\text{Bribe}^{CM}},p^{CM}p^{BP},q)$
@@
    • Repeats the same procedure for the includers with order $2,\dots,j-1$.